OpenStudy (anonymous):

Find the sum of the first ten terms of the sequence -4, 1, 6, ........41 in the tenth term. The answer is 185. I need help finding the rate, and the set up of the equation, I can't seem to figure it out.

OpenStudy (anonymous):

This is an Arithmatic progression where each term is 5 more than the previous term

OpenStudy (anonymous):

In such a case, the value of the nth terms is given by the formula \[s _{n} = n/2 (a + l)\] where sn stands for sum a = first term which is -4 in this case l = last term which is 41 in this case n = no of term u want to find which is 10 in yr case

OpenStudy (anonymous):

so putting in the values u get \[s _{10} = 10/2 (-4 + 41) = 5 x 37 = 185\]

OpenStudy (anonymous):

Does this work for you????

OpenStudy (anonymous):

OMG!!!! YEEEEESSSSS!!! That is the most elaborate answer I have ever gotten EVER!!! Thank you sooooo much!!!!!!!!!

OpenStudy (anonymous):

I'm a teacher by profession, so I try to giv ans which r crystal clear.......... Medal for me !!! :)

OpenStudy (anonymous):

Well I wish you were my teacher!! I have been having such a hard time understanding math this semester.... Thank you so much, all I needed was for the break down!! Thank you thank you thank you!!!

OpenStudy (anonymous):

Good Job Harkirat...that's the way I would have explained it

OpenStudy (anonymous):

Exactly edbraunh!!

OpenStudy (anonymous):

thanks edbraunch u r kind

OpenStudy (anonymous):

I am frm India and I am already helping a kid in US thru e-tuition so I know how to giv solution so tht the other party understands

OpenStudy (anonymous):

You are a great teacher I thank you again to very much!!! Your one of a kind =)))

OpenStudy (anonymous):

U r most welcome !! I suggest you collect all the formulas related to Arithmetic progression, understand the use of each one and memorize them. Then u will b on top of yr sums (and hopefully yr class too!!)

OpenStudy (anonymous):

Thank you!! And I will